A rectangle has a length of 38 meters less than 8 times its width. If the area of the rectangle is 4050 square meters, find the length of the rectangle.
Answer by TimothyLamb(4379)

L = 8w - 38
Lw = 4050
---
(8w - 38)w = 4050
8ww - 38w - 4050 = 0
---
the above quadratic equation is in standard form, with a=8, b=-38, and c=-4050
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
8 -38 -4050
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
w = 25
w = 25
---
the negative root doesn't fit the problem statement, so use the positive root
---
w = 25 meters
---
answer:
L = 4050/w = 4050/25 = 162 meters
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php